OpenStudy (anonymous):

find dy/dx y=cos^-1[(1-x^2)/(1+x^2)]

OpenStudy (anonymous):

\[y = \cos ^{-1}(\frac{ 1-x ^{2} }{ 1+x ^{2} })\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

step1; find the derivative of cos inverse

OpenStudy (anonymous):

step2; use chain rule for the function inside, you would need to chain rule

OpenStudy (anonymous):

and quotient rule

OpenStudy (anonymous):

okay , what have told me , i have done all that my answer coming is 2/(1+x^2)

OpenStudy (anonymous):

so, is my anwer correct

OpenStudy (anonymous):

where did cos inverse disappear? It will still be there

OpenStudy (anonymous):

so first you find cos inverse derivative, then multiply it with the chain rule of the function inside cos inverse

OpenStudy (anonymous):

your ans would be derivative of cos inverse x chain rule

OpenStudy (anonymous):

k what i did is i write cos y =(1-x^2)/(1+x^2) then i solve it

OpenStudy (anonymous):

cosy =( 1-tan^2y/2)/(1+tan^2y/2) ( 1-tan^2y/2)/(1+tan^2y/2)=( 1-x^2/2)/(1+x^2/2)

OpenStudy (anonymous):

on comparing, we get, tan(y/2)= x then i differentiate wrt x

OpenStudy (anonymous):

then i get dy/dx= 2 /sec^2(y/2) dy/dx=2/[1+tan^2(y/2)] dy/dx=2/[1+x^2]

OpenStudy (anonymous):

this is my answer

OpenStudy (anonymous):

@stamp what r u trying to say

OpenStudy (anonymous):

@Kanwar245 plse tell me is my answer right

OpenStudy (anonymous):

u mean \[\arccos *\frac{ 1-x^2 }{ 1+x^2 }\]

OpenStudy (anonymous):

it is coming to be \[\frac{ (x^2+1)(-2x \cos^{-1}x -\sqrt{(1-x^2)}-2x(1-x^2)\cos^{-1} x }{ (x^2+1)^2 }\]